Car agency A will rent a compact car for $40 per day and an additional charge of $0.20 per mile. Car rental agency B will charge only $0.16 per mile but charges $51 per day. Adam wanted to rent a car for four days, how many miles would Adam have to drive to make car rental agency B a better bargain?
Answer by checkley75(3666)

4*40+.20x=4*51+.16x
160+.20x=204+.16x
.20x-.16x=204-160
.04x=44
x=44/.04
x=1,100 miles driven makes the costs for the 2 plans equal.
Proof:
4*40+.20*1,100=4*51+.16*1,100
160+220=204+176
380=380